OpenStudy (anonymous):

Please help!!!!!! lim as x approaches 0 of (2(sin3x)^3/-x^3)

OpenStudy (rajee_sam):

\[\lim_{x \rightarrow 0}\frac{ 2 Sin ^{3} 3x }{ -x ^{3} }\] is this your problem?

OpenStudy (anonymous):

Yes!

OpenStudy (rajee_sam):

do you know L'Hopital's rule?

OpenStudy (anonymous):

no I do not

OpenStudy (rajee_sam):

if you apply the limit as it is you will get a form 0/0 . So whenever you face such a situation where your limit turns to the form 0/0, or infinity / infinity, -infinity/infinity, -inifinty/-infinity, you an apply L'Hopitals rule.

OpenStudy (rajee_sam):

For using L'Hopital's rule you need to take the derivative of the numerator separately and the derivative of the denominator separately. An then apply the limit again. If you face the same situation as earlier, You do the derivative one more time. Then apply the limit. You keep on taking the derivative until you reach a determinate limit.

OpenStudy (rajee_sam):

here if you apply limit 0 the whole thing becomes 0/0 Now take the derivative of the numerator ( Apply Chain rule) Seperately. Take the derivative of denominator separately.

OpenStudy (anonymous):

So I take the derivative of each portion separately until I get a constant in the denominator?

OpenStudy (rajee_sam):

yes

OpenStudy (rajee_sam):

in this case

OpenStudy (rajee_sam):

but if you have different thing in the denominator say you had Sin x. If you differentiate that it becomes Cosx and if you apply 0 its value is 1.

OpenStudy (anonymous):

Thank you! I believe you get -- the third derivative of f =\[\frac{ -12\cos(3x) }{ -6 }\]

OpenStudy (anonymous):

do I solve from there by plugging in 0 for x?

OpenStudy (anonymous):

in the end you get 2 -- yes?

OpenStudy (rajee_sam):

for differentiating the numerator you apply chain rule uv

OpenStudy (rajee_sam):

how do you get just -12 cos 3x

OpenStudy (anonymous):

yes but by applying the chain rule - you just multiple the numerator by 1

OpenStudy (rajee_sam):

\[2 \sin ^{3} 3x\] that is the numerator right?

OpenStudy (anonymous):

the first derivate of the numerator is: \[6 \cos ^{2} (3x)\]

OpenStudy (anonymous):

yes

OpenStudy (rajee_sam):

first derivative is 6Sin²3x . Cos 3x . 3 = 18 Sin²3x .Cos 3x

OpenStudy (rajee_sam):

Sin³ 3x = 3. Sin² 3x . Cos 3x. 3 = 9 Sin² 3x. Cos 3x you already had a 2 in the front

OpenStudy (rajee_sam):

now you have to do uv method

OpenStudy (rajee_sam):

2nd derivative 18 sin² 3x . Cos 3x = 18[ 2 Sin 3x Cos 3x. 3 . Cos 3x + Sin² 3x . (- Sin 3x) . 3 ] = 18 [ 6 . Sin 3x. Cos² 3x - 3 Sin³ 3x] = 108 Sin 3x. Cos² 3x - 54 Sin³ 3x

OpenStudy (rajee_sam):

3rd derivative 108 [ Cos 3x. 3 . Cos² 3x + Sin 3x . 2 Cos 3x . (-sin 3x) . 3] - 54[ 3Sin²3x . Cos 3x . 3] = 108 [ 3 Cos³ 3x - 6 Sin² 3x . Cos 3x ] - 486 Sin² 3x . Cos 3x = 324 Cos³ 3x - 648 Sin² 3x . Cos 3x - 486 Sin² 3x . Cos 3x = 324 Cos³ 3x - 1134 Sin² x . Cos 3x

OpenStudy (rajee_sam):

now if you apply the limit the second term becomes 0. and the first term becomes 324 the denominator you have -6 so your answer will be 324 / -6 = -54